# Toward TransfORmers: Revolutionizing the Solution of Mixed Integer Programs with Transformers

### v1

@@ Abstract
In this study, we introduce an innovative deep learning framework that employs a transformer model to address the challenges of mixed-integer programs, specifically focusing on the Capacitated Lot Sizing Problem (CLSP). Our approach, to our knowledge, is the first to utilize transformers to predict the binary variables of a mixed-integer programming (MIP) problem. Specifically, our approach harnesses the encoder decoder transformer’s ability to process sequential data, making it well-suited for predicting binary variables indicating production setup decisions in each period of the CLSP. This problem is inherently dynamic, and we need to handle sequential decision making under constraints. We present an efficient algorithm in which CLSP solutions are learned through a transformer neural network. The proposed post-processed transformer algorithm surpasses the state-of-the-art solver, CPLEX and Long Short-Term Memory (LSTM) in solution time, optimal gap, and percent infeasibility over 240K benchmark CLSP instances tested. After the ML model is trained, conducting inference on the model, including post-processing, reduces the MIP into a linear program (LP). This transforms the ML-based algorithm, combined with an LP solver, into a polynomial-time approximation algorithm to solve a well-known NP-Hard problem, with almost perfect solution quality.
@@ 2.1 CLSP: Mixed Integer Program Formulation
The Capacitated Lot Size Problem (CLSP) is an optimization challenge that involves determining optimal production quantities for multiple items over a series of time periods, subject to production capacity constraints, with the aim of minimizing total costs, including setup and inventory holding costs. Let $T$ represent the number of periods considered within the planning horizon, where $t\in\{1,2,\ldots,T\}$ is the index that represents each period within the planning horizon. Let $d_{t}$, $p_{t}$, $f_{t}$, $h_{t}$, and $c_{t}$ represent the expected demand, the unit production cost, the setup cost, the unit inventory holding cost, and the production capacity in period $t$, respectively. Given that $x_{t}$, $s_{t}$, and $y_{t}$ are the variables that define the units produced, the ending inventory, and the binary production setup variable that is equal to 1 if there is production in period $t$, and 0 otherwise, in period $t$, the CLSP MIP formulation is presented as follows:
$s\min$ $\sum_{t=1}^{T}(p_{t}x_{t}+f_{t}y_{t}+h_{t}s_{t})$ (Objective Function) (1)
@@ 2.3 Data and Preprocessing
The model was trained with synthetically generated benchmark data following the CLSP instance generation schema of Atamtürk and Munoz (2004): capacity-to-demand ratios $c\in\{3,5,8\}$, setup-to-holding cost ratios $f\in\{1000,10000\}$ and the number of periods $T\in\{90\}$. The parameters regarding demand $d_{t}$, unit production cost $p_{t}$, production capacity $c_{t}$, and setup cost $f_{t}$ are generated from integer uniform distribution with the ranges, $d_{t}\in[1,600]$, $p_{t}\in[1,5]$, $c_{t}\in[0.7c,1.1c]$, $f_{t}\in[0.9f,1.1f]$, where $c$ and $f$ are defined in parameters. The inventory holding cost $h_{t}$ is set to one. The CLSP instances were solved using CPLEX 22.1.0 and optimal solutions, objective function value, and solve times were stored in our dataset for validation.
The data set comprised 1,200,000 synthetic randomized instances of CLSP, generated and solved to create a complete training and testing dataset (Yilmaz and Büyüktahtakın, 2023). These synthetic data provide a robust platform for the development and validation of models. 20% of the data were used for validation and testing, the remaining data 80 % making up the training data. Each six of the $c$ and $f$ combinations have 40,000 instances for validation and testing. Pre-processing for the transformer involved batch normalization and log-scaling of the input data. Additionally, a vector embedding scheme was employed for numerical values similar to Vaswani et al. (2017)’s embedding scheme, but deviating from traditional word-focused embeddings, to better suit the numerical nature of CLSP data. The model presented was trained for .95 GPU hours.
@@ 2.4 Post-Processing Transformer Predictions
The predicted $y_{t}$-variables are fixed in the CLSP MIP model (1)–(5), reducing the problem into an LP. Then the resulting LP is solved by CPLEX 22.1.0. Our prediction results were accurate, with the exception of the prediction of the variable $y_{t}$ at time $T$ in 2% of the test cases, as shown in Table 2, where the column ‘Inf’ gives the percentage of instances that are infeasibly predicted for each of the six combinations of $c$ and $f$.
@@
Table 2: Model infeasibility rate with no post-processing
To address the imperfect predictions in the final period, we present a heuristic that runs two potential scenarios through CPLEX in parallel and selects only the feasible outcome, as demonstrated in Figure 1. This approach empirically results in only optimal feasible solutions as output while significantly improving the solve time as discussed in detail in the next section.
@@ 3.1 Model Performance
The transformer model demonstrates exceptional performance in predicting the binary variables of the CLSP, as detailed in Table 3. The ‘TS Solve Time’ gives the combined CPU and GPU seconds to solve the CLSP instance with the integrated transformer framework (computational experiments were performed with an A100 GPU and 2 CPU cores), the percentage of infeasibility (Inf
@@ C $f$ TS Solve Time Optgap(%) Inf (%)
Table 3: Summary of Average Computational Results after Post-Processing.
Noticeably, the model performs better on larger $f$ values (Figure 2). This is probably a function of the frequency of production in the last period for each configuration in the data set.
@@ 3.2 Comparative Analysis
c
f TimeCPX
TimeML
Timegain(%) Inf(%) Optgap(%)
LSTM
TS LSTM TS LSTM TS LSTM TS
3
1000 0.107
0.02
0.003 81.31 97.20 0.00 0.00 0.59 0.00
5
1000 0.07
0.02
0.003 71.43 95.71 1.17 0.00 0.16 0.00
8
1000 0.06
0.02
0.003 66.67 95.00 0.00 0.00 0.05 0.00
3
10000 0.79
0.02
0.003 97.46 99.62 67.70 0.00 0.80 0.00
5
10000 0.65
0.02
0.003 96.92 99.54 42.00 0.00 1.30 0.00
8
10000 0.33
0.02
0.003 93.93 99.09 22.60 0.00 3.18 0.00
Average:
0.3342
0.02
0.003 94.02 99.10 22.245 0.00 1.013 0.00
@@
The model outperforms CPLEX in terms of solve time with an improvement of more than 99% on average. The model is also superior to the LSTM in terms of solution quality, inference time, and training time.
@@ 3.3 Significance of Research Results
As presented in the former section, the proposed post-processed transformer algorithm surpasses the state-of-the-art solver, CPLEX, and Long Short-Term Memory (LSTM) in solution time, optimal gap, and percent infeasibility over 240K benchmark CLSP instances tested. The results underscore the potential of transformer models to effectively solve mixed-integer programs, particularly in cases where rapid and repeated solving is crucial. Despite the fact that transformers have fewer inductive biases for sequential problems compared to LSTM architectures, they showed an exceptional ability to learn binary variables in MIPs. These findings position transformer models as a state-of-the-art approach to deep learning methods for MIPs, far surpassing traditional heuristics.
@@ 4 Discussion and Future Work
We present an efficient algorithm in which CLSP solutions are learned through a transformer neural network. The proposed post-processed transformer algorithm achieves state-of-the-art performance in solution time, optimal gap, and percent infeasibility. After training the model, running inference with post-processing on in-distribution data relaxes the problem to a Linear Programming (LP) instance, solvable in polynomial time. This transformation effectively converts the learned algorithm into a polynomial-time solution for a well-known NP-Hard problem.
Furthermore, the model finds the optimal solution in 100% of the test cases when simple post-processing is performed. Figure 2 shows no noticeable patterns in attention weights, confirming what researchers have known for 50 years, that there is no simple solution to hard NP problems (Johnson, 2012). Rather, the performance of the transformer model on the CLSP with these attention heads can be attributed to the transformer’s ability to ’understand’ very complex relationships.
Figure 2: Transformer Attention heatmap for 2 Attention head transformer
@@ 4.2 Limitations
Compared to traditional heuristics and other machine learning approaches, our transformer-based method demonstrates superior or equivalent performance, with the added benefit of faster training times. However, it is important to note that transformers are data intensive and require a deep understanding of deep learning for effective training and tuning. This limitation suggests that, while transformers offer considerable advantages, they are not a universal solution for all combinatorial optimization challenges. The finicky nature of these models could affect their broader applicability, as proper training and tuning are critical for optimal performance. This limitation underscores the need for expertise in deep learning to harness the full potential of transformer models in operations research applications. Additionally, the inductive biases of other architectures may be preferential for some MIPs, such as knapsack or traveling salesman problems. The model fails to predict the last period of the CLSP in 2% of instances. Although the exact dynamics responsible for this behavior is unknown, positional bias is likely to be the culprit (Mehdi B. Amor, 2023).

### v3

@@ Abstract
In this study, we introduce an innovative deep learning framework that employs a transformer model to address the challenges of mixed-integer programs, specifically focusing on the Capacitated Lot Sizing Problem (CLSP). Our approach, to our knowledge, is the first to utilize transformers to predict the binary variables of a mixed-integer programming (MIP) problem. Specifically, our approach harnesses the encoder decoder transformer’s ability to process sequential data, making it well-suited for predicting binary variables indicating production setup decisions in each period of the CLSP. This problem is inherently dynamic, and we need to handle sequential decision making under constraints. We present an efficient algorithm in which CLSP solutions are learned through a transformer neural network. The proposed post-processed transformer algorithm surpasses the state-of-the-art solver, CPLEX and Long Short-Term Memory (LSTM) in solution time, optimal gap, and percent infeasibility over 240K benchmark CLSP instances tested. After the ML model is trained, conducting inference on the model, reduces the MIP into a linear program (LP). This transforms the ML-based algorithm, combined with an LP solver, into a polynomial-time approximation algorithm to solve a well-known NP-Hard problem, with almost perfect solution quality.
@@ 2.1 CLSP: Mixed Integer Program Formulation
The Capacitated Lot Size Problem (CLSP) is an optimization challenge that involves determining optimal production quantities for items over a series of time periods, subject to production capacity constraints, with the aim of minimizing total costs, including setup and inventory holding costs. Let $T$ represent the number of periods considered within the planning horizon, where $t\in\{1,2,\ldots,T\}$ is the index that represents each period within the planning horizon. Let $d_{t}$, $p_{t}$, $f_{t}$, $h_{t}$, and $c_{t}$ represent the expected demand, the unit production cost, the setup cost, the unit inventory holding cost, and the production capacity in period $t$, respectively. Given that $x_{t}$, $s_{t}$, and $y_{t}$ are the variables that define the units produced, the ending inventory, and the binary production setup variable that is equal to 1 if there is production in period $t$, and 0 otherwise, in period $t$, the CLSP MIP formulation is presented as follows:
$\min$ $\sum_{t=1}^{T}(p_{t}x_{t}+f_{t}y_{t}+h_{t}s_{t})$ (Objective Function) (1)
@@ 2.3 Data and Preprocessing
The model was trained with synthetically generated benchmark data following the CLSP instance generation schema of Atamtürk and Munoz (2004): capacity-to-demand ratios $c\in\{3,5,8\}$, setup-to-holding cost ratios $f\in\{1000,10000\}$ and the number of periods $T\in\{90\}$. The parameters regarding demand $d_{t}$, unit production cost $p_{t}$, production capacity $c_{t}$, and setup cost $f_{t}$ are generated from integer uniform distribution with the ranges, $d_{t}\in[1,600]$, $p_{t}\in[1,5]$, $c_{t}\in[0.7c,1.1c]$, $f_{t}\in[0.9f,1.1f]$, where $c$ and $f$ are defined in parameters. The inventory holding cost $h_{t}$ is set to one. CLSP instances were solved using CPLEX 22.1.0 and optimal solutions, objective function value, and solve times were stored in our dataset for validation.
The data set comprised 1,200,000 synthetic randomized instances of CLSP, generated and solved to create a complete training and testing dataset (Yilmaz and Büyüktahtakın, 2023). These synthetic data provide a robust platform for the development and validation of models. 20% of the data were used for validation and testing, the remaining data 80 % making up the training data. Each six of the $c$ and $f$ combinations have 40,000 instances for validation and testing. Pre-processing for the transformer involved batch normalization and log-scaling of the input data. Additionally, a vector embedding scheme was employed for numerical values similar to Vaswani et al. (2017)’s embedding scheme, but deviating from traditional word-focused embeddings, to better suit the numerical nature of CLSP data. The model presented was trained for 0.95 GPU hours.
@@ 2.4 Post-Processing Transformer Predictions
The predicted $y_{t}$-variables are fixed in the CLSP MIP model (1)–(5), reducing the problem into an LP. Then the resulting LP is solved by CPLEX 22.1.0. Our initial model’s prediction results were accurate, except for the prediction of the variable $y_{t}$ at time $T$ in 2% of the test cases, as shown in Table 2, where the column ‘Inf’ gives the percentage of instances that are infeasibly predicted for each of the six combinations of $c$ and $f$.
@@
Table 2: Model infeasibility rate with no post-processing or $\langle EOS\rangle$ token.
To address the imperfect predictions in the final period, we present a heuristic that runs two potential scenarios through CPLEX in parallel and selects only the feasible outcome, as demonstrated in Figure 1. This approach empirically results in only optimal feasible solutions as output while significantly improving the solve time as discussed in detail in the next section. We also found that the use of an End of Sequence $\langle EOS\rangle$ token eliminates the infeasiblity while incurring very marginally higher training time and lower inference time compared to our post-processing techniques.
@@ 3.1 Model Performance
The transformers model (TS) demonstrates exceptional performance in predicting the binary variables of the CLSP, as detailed in Table 3. The TS Solve Time’ gives the combined CPU and GPU seconds to solve the CLSP instance with the integrated transformer framework (computational experiments were performed with an A100 GPU and 2 CPU cores while CLPEX solve times used 20 cores). The percentage of infeasibility (Inf %) represents the proportion of infeasible solutions generated by the transformer model, while the optimality gap (Optgap %) illustrates the percentage difference between the optimal solution obtained by CPLEX and the solution achieved by integrating the predictions of the transformer model with CPLEX. Our integrated transformer framework achieves an infeasibility percentage (Inf %) and an optimality gap (Optgap %) of 0%, while significantly reducing CPLEX solve times by more than 99%.
@@ C $f$ TS Solve Time Optgap(%) Inf (%)
Table 3: Summary of Average Computational Results with Post-Processing or $\langle EOS\rangle$ token.
Noticeably, the initial model performs better with larger $f$ values (Figure 2). This is probably a function of the frequency of production in the last period for each configuration in the data set.
@@ 3.2 Comparative Analysis
{tblr}
width = colspec = Q[27]Q[50]Q[20]Q[30]Q[67]Q[67]Q[67]Q[69]Q[69]Q[67]Q[67]Q[67]Q[67], cell13 = c=10.1, cell15 = c=20.1, cell17 = c=20.1, cell19 = c=20.1, cell111 = c=20.1, cell113 = c=20.1, cell31 = r, cell32 = r, cell41 = r, cell42 = r, cell51 = r, cell52 = r, cell61 = r, cell62 = r, cell71 = r, cell72 = r, cell81 = r, cell82 = r, hline1,9 = -0.08em, hline2 = -, hline10,9 = -0.08em, c & $f$ TimeCPX TimeML Timegain(%) Inf(%) Optgap(%)
LSTM TS LSTM TS LSTM TS LSTM TS
3 1000 0.107 0.02 0.003 81.31 97.20 0.00 0.00 0.59 0.00
5 1000 0.07 0.02 0.003 71.43 95.71 1.17 0.00 0.16 0.00
8 1000 0.06 0.02 0.003 66.67 95.00 0.00 0.00 0.05 0.00
3 10000 0.79 0.02 0.003 97.46 99.62 67.70 0.00 0.80 0.00
5 10000 0.65 0.02 0.003 96.92 99.54 42.00 0.00 1.30 0.00
8 10000 0.33 0.02 0.003 93.93 99.09 22.60 0.00 3.18 0.00
Average: 0.334 0.020.003 94.02 99.10 22.245 0.00 1.013 0.00
@@
$c$ $f$ TimeGRB Time⟨EOS⟩ Timegain(%) Inf(%) Optgap(%)
3 1000 0.06 0.003 95.00 0.00 0.00
5 1000 0.13 0.003 97.69 0.00 0.00
8 1000 0.10 0.003 97.00 0.00 0.00
3 10000 0.07 0.003 95.71 0.00 0.00
5 10000 0.12 0.003 97.50 0.00 0.00
8 10000 0.10 0.003 99.09 0.00 0.00
Average 0.096 0.003 96.875 0.00 0.00
Table 5: Performance comparison using Transformer (TS) with $T=90$ period instances using an ⟨EOS⟩token. Results are using 10k instances per each $c$ and $f$ configuration. TimeGRB is solve time using Gurobi with the facet-defining $(\ell,S)$ inequalities proposed by Barany et al. (1984).
The model outperforms CPLEX in terms of solve time with an improvement of more than 99% on average and outperforms Gurobi with (l,S) inequalities by 96.9%. The model is also superior to the LSTM in terms of solution quality, inference time, and training time.
@@ 3.3 Significance of Research Results
As presented in the former section, the proposed post-processed transformer algorithm and transformer with $\langle EOS\rangle$ token surpasses the state-of-the-art solvers, CPLEX and Gurobi, and Long Short-Term Memory (LSTM) in solution time, optimal gap, and percent infeasibility over 240K benchmark CLSP instances tested. The results underscore the potential of transformer models to effectively solve mixed-integer programs, particularly in cases where rapid and repeated solving is crucial. Despite the fact that transformers have fewer inductive biases for sequential problems compared to LSTM architectures, they showed an exceptional ability to learn binary variables in MIPs. These findings position transformer models as a state-of-the-art approach to deep learning methods for MIPs, far surpassing traditional heuristics.
@@ 4 Discussion and Future Work
We present an efficient algorithm in which CLSP solutions are learned through a transformer neural network. The proposed post-processed transformer algorithm achieves state-of-the-art performance in solution time, optimal gap, and percent infeasibility. After training the model, running inference with post-processing on in-distribution data relaxes the problem to a Linear Programming (LP) instance, solvable in polynomial time. This transformation effectively converts the learned algorithm into a polynomial-time solution for a well-known NP-Hard problem. The model finds the optimal solution in 100% of the test cases when trained with an $\langle EOS\rangle$ token or simple post-processing is performed.
@@ 4.2 Limitations
Compared to traditional heuristics and other machine learning approaches, our transformer-based method demonstrates superior or equivalent performance, with the added benefit of faster training times. However, it is important to note that transformers are data intensive and require an in-depth understanding of deep learning for effective training and tuning. This limitation suggests that, while transformers offer considerable advantages, they are not a universal solution for all combinatorial optimization challenges. The finicky nature of these models could affect their broader applicability, as proper training and tuning are critical for optimal performance. This limitation underscores the need for expertise in deep learning to harness the full potential of transformer models in operations research applications. Additionally, the inductive biases of other architectures may be preferential for some MIPs, such as knapsack or traveling salesman problems. The base model fails to predict the last period of the CLSP in 2% of instances without end of sequence token. Although the exact dynamics responsible for this behavior is unknown, positional bias is likely to be the culprit (Mehdi B. Amor, 2023).